OpenStudy (zolock):

Simplify Square rot of 1008. a. 18*square root of 3 b. 12*square root of 7 c. 12*square root of 6 d. 11*square root of 7

OpenStudy (anonymous):

B

OpenStudy (jdoe0001):

\(\bf 1008\implies 2\cdot 2\cdot 2\cdot 2\cdot 3\cdot 3\cdot 7\implies 2^2\cdot 2^2\cdot 3^2\cdot 7 \\ \quad \\ \Large \sqrt{1008}\implies \sqrt[{\color{red}{ 2}}]{1008}\implies \sqrt[{\color{red}{ 2}}]{2^{\color{red}{ 2}}\cdot 2^{\color{red}{ 2}}\cdot 3^{\color{red}{ 2}}\cdot 7}\)